CASSCF CALCULATIONS VIA VARIATIONAL TWO-ELECTRON REDUCED-DENSITY-MATRIX THE-ORY

\underline{GERGELY GIDOFALVI}, Chemical Sciences and Engineering Division, Argonne National Laboratory, Argonne, IL 60439; DAVID A. MAZZIOTTI, Department of Chemistry and The James Franck Institute, The University of Chicago, Chicago, IL 60637.

Variational reduced-density-matrix theory allows the computation of ground-state energies without the explicit construction of the N -electron wavefunction, and, with the incorporation of orbital rotations, may be employed in CASSCF calculations. The implementation and application of the method will be presented^{a}.

aG. Gidofalvi and D. A. Mazziotti, J. Chem. Phys. 127, 244105 (2007)